# Large Language Models as Span Annotators

### v1

@@ 6 Discussion
Our results paint a complex picture in this regard. The IAA between LLMs and human annotators is generally moderate. However, we also need to take into account a realistic upper bound: the average IAA between human annotators themselves. In that case, the strongest LLMs have already reached the level of human annotators in D2T-Eval and are not far behind in MT-Eval and Propaganda. As we also showed, not all human annotations are “gold”, as even qualified crowdworkers make similar amounts of errors as LLMs. If we also factor in financial aspects, LLMs provide better flexibility, shorter response times, and much lower costs. In light of this, we argue that the benefits of automating span annotation with LLMs already outweigh its issues. Moreover, LLM-based span annotation will benefit from future increases in LLM capabilities.
@@
To instantiate an LLM as a span annotator, it is beneficial to provide detailed guidelines describing how to handle all edge cases. In contrast, we note that providing specific examples (i.e., a few-shot setup) does not bring consistent improvements due to the length and complexity of typical examples for this task. When using LLMs in a new setup, we recommend validating model’s annotations against examples hand-annotated by experts on a sample of the data. As for choosing the model, we recommend using a reasoning model (as opposed to an equivalent model without reasoning), since the reasoning models provide more reliable annotations, although at the cost of slightly higher response times and token count.
@@ 7 Conclusion
We demonstrated that LLMs can serve as span annotators for three span annotation tasks: evaluation of data-to-text generation, evaluation of machine translation, and propaganda detection in human-written texts. Our experiments show that LLMs achieve moderate agreement with skilled human annotators. The models perform best in D2T-Eval, where they are comparable to qualified crowdworkers. Reasoning models consistently outperform their instruction-tuned counterparts, delivering more accurate annotations and providing more valid explanations for their decisions. Overall, automating span annotation with LLMs offers a promising alternative to human annotation, opening the way for more fine-grained NLG evaluation methods.
@@ Acknowledgments
Our human annotators were hired over Prolific and paid the platform-recommended wage of 9 GBP/hour (adjusted to slightly higher rates to account for real annotation times). The annotators were preselected based on their primary language (English). All annotators were shown detailed instructions and explanation of the data types, data sources, and the purpose of the research (see Appendix D.2 for details). The domains were selected so that they do not contain any sensitive or potentially offensive content. We do not collect any demographic data about the participants.
@@ Appendix A Implementation Details
A.1 Costs and Time Requirements
We estimate the cost and time requirements for a representative subset of annotators on D2T-Eval in Table 4.
Metric crowdw. Llama 3.3 DeepSeek-R1 Claude 3.7 Sonnet o3-mini
Cost per 1k outputs (\$) 500 - - 10.5 3.6
Time per output (s) 129.1 21.6 227.5 9.0 21.8
Table 4: Estimate of costs and time requirements for different kinds of annotators on D2T-Eval: crowdworkers on Prolific, open models (Llama 3.3, DeepSeek-R1), and two out of the proprietary models (Claude 3.7 Sonnet, and o3-mini).
Note that we do not estimate the costs of the models we run locally. In practice, the costs of infrastructure acquisition and inference (GPUs, electricity) needs to be taken into account. The time estimates for the open models were computed on the infrastructure we describe in Section A.2.
A.2 Open Models
@@
For better reproducibility, we set the seed to 42 and temperature to 0 for the local models. We do not use these parameters for proprietary models, as these parameters are generally not supported.
@@ A.3 Proprietary Models
Appendix B Extracting Annotations from LLM Outputs
B.1 Tagging the Input Text
To the best of our knowledge, there is no established way of tagging the input text with decoder-only models. In the following, we discuss some of the existing approaches along with their pros and cons.
  1. 1.
    Labeling the text with special tokens Wang et al. (2023); Treviso et al. (2024): This approach instructs the LLM to generate the full copy of the input text. Spans are denoted using special XML-like tokens added to this copy. The approach is unambiguous in case the model copies the text correctly. However, it relies on the ability of the LLM to produce an exact copy of the (potentially very long) input text, and offers no fallback if the model fails to do that.
  2. 2.
    Listing start and end indices of spans Hasanain et al. (2024); Ramponi et al. (2025): This approach asks the LLM to generate numerical indices of the characters or tokens in the span. The approach would also make the output unambiguous, but it was discouraged by Hasanain et al., who notes that “[GPT-4] frequently generated indices not matching the corresponding spans in a paragraph”. Similarly, when models were asked to produce the output in CoNLL format Sang & Meulder (2003), “the models struggled to provide consistent outputs” Ramponi et al. (2025). A potential solution, providing the identifier for each token in the input text Ramponi et al. (2025), could be a confounding factor as it modifies the input text.
  3. 3.
    Listing textual content of spans Kasner & Dušek (2024); Kocmi & Federmann (2023); Zouhar et al. (2024): This approach asks the LLM to produce the textual content of the labeled spans along with their labels. The start and end indices are then found heuristically by case-insensitive string matching on the input text. This approach is more flexible, as the model does not need to copy the entire text, although it still relies on the ability of the models to literally copy text snippets. It can also misalign the model output by assigning the annotation to the first match in case there are multiple matches.
Of the available options, we select listing textual content of spans as our method of choice. Although we are aware of its disadvantages, our preliminary experiments showed that the failures are infrequent: the models are generally able to copy text snippets literally and string matching collisions are rare. This approach is also the most robust, as failing to match a single annotation does not influence other matches.
    "annotations": [{
      "reason": [REASON],
      "text": [SPAN_TEXT],
      "type": [CATEGORY_IDX]
Figure 5: Annotation output format for LLMs.
B.2 Structured Output
Having a high-level idea of the output format, we need to determine the exact format required from the model. Existing works collect model annotations using unstructured lists Kocmi & Federmann (2023); Zouhar et al. (2024) and structured output in the JSON format Kasner & Dušek (2024).
Following our preliminary experiments, we decide to closely follow the setup of Kasner & Dušek (2024), using structured decoding with a fixed JSON scheme as we describe in Section 3.2 (see Figure 5). Our rationale is that JSON outputs are easier and more robust to parse, while being a common format that is natural for the LLMs. Structured output is now also getting available in LLM frameworks. [^12] See, e.g., https://ollama.com/blog/structured-outputs.
Can structured output hurt model performance?
Some authors voice concerns about the diminished performance of LLMs when using JSON output Tam et al. (2024); Beurer-Kellner et al. (2024). However, reactions to Tam et al. Kurt (2024); Castillo (2024) suggest that these issues do not manifest if several factors—such as describing the JSON format in the prompt—are controlled for. We leave experiments on this issue as a separate topic for future work.
Parsing the output with reasoning chains
@@ B.2 Structured Output
At the time we performed our experiments, some of the reasoning models either did not support structured output (Gemini 2.0 Flash Thinking) or it would cancel out their ability to perform reasoning chains (ollama models [^13] https://github.com/ollama/ollama/issues/8529). Therefore, we use the following heuristic: we retrieve the raw answer from the model, strip any parts within the <think></think> tags (if present), and consider the latest valid top-level JSON object as the model’s response. This approach turns out to be able to robustly parse model responses containing reasoning chains with minimum failures.
Appendix C Discussion on Data Contamination
It is common practice not to disclose the data used for LLM training. That applies not only to the prorietary models, but also to the majority of the open models. As discussed in several recent works Balloccu et al. (2024); Dong et al. (2024); Jiang et al. (2024b), it is mostly certain that the LLMs were trained on some test sets of benchmarks that were available before their knowledge cut-off date. Memorization of these test sets may then artificially inflate the model performance on these benchmarks.
In the context of D2T generation, Kasner & Dušek (2024) suggests to download a new set of structured data from publicly available sources – e.g., downloading up-to-date weather data for weather forecast generation. We use the method for D2T-Eval, where it creates only little overhead.
For the other tasks, we are not able to follow the same process due to the costs of data collection. Nevertheless, we believe that it does not influence the validity of our results:
  1. 1.
  2. 2.
    For Propaganda, the dataset is in a format that cannot be easily memorized. The model would need to match the word-level annotations for token ids with the text, which (as we also discuss in Section B.2) is not typically within model capabilities.
  3. 3.
    Lastly, D2T-Eval serves as a control task. If the model would underperform on D2T-Eval, but perform well on other tasks where the datasets were known, then we could suspect that the models memorized the data. In practice, we see the opposite: the models have the lowest performance on Propaganda which is the only task with a publicly available dataset.
Appendix D D2T-Eval Details
@@
In this section, we provide more details on the process of collecting data (Section D.1) and annotations (Section D.2) for D2T-Eval.
D.1 Collecting Data
@@
Our goal was to collect a novel dataset for D2T-Eval that is not subject to data contamination (cf. Appendix C). To this end, we follow the approach of Kasner & Dušek (2024) to collect structured data from publicly available APIs.
@@
The Quintd tool Kasner & Dušek (2024) enables downloading data for five domains: openweather, gsmarena, ice_hockey, our_world_in_data, and wikidata. We selected openweather and gsmarena as the most suitable for the data collection process. Additionally, we collect the new football domain from RapidAPI - API-Football. [^14] Our investigation showed that its structured data contain more detailed information about the game than ice_hockey and the sport itself is also more generally well-known.
@@
We leverage LLMs to get output texts for the dataset. Specifically, we prompt two state-of-the-art models, Llama 3.3 Grattafiori et al. (2024) and GPT-4o Hurst et al. (2024), and two smaller models, Gemma 2 Team et al. (2024) and Phi-3.5 Abdin et al. (2024) (cf. Section A.2). We prompt the models in a zero-shot setting, asking them to generate a summary of the given data using approximately five sentences. We provide the prompts in Appendix F.
@@
We point out that two of the four models that we use to collect annotations (Llama 3.3 and GPT-4o) are also later used to collect annotations. The fact that the models annotate their own outputs can reveal the self-preference bias Zheng et al. (2023); Wataoka et al. (2024). Since our main point is not to rank the quality of generated outputs, we do not consider it a flaw, but rather a suitable ground for future investigation. We hypothesize that since span annotations are more grounded than direct assessment (and not directly ranking models), the self-preference bias will not be as prominent here.
D.2 Collecting annotations
Here we describe our process of writing annotation guidelines and selecting crowdworkers in order to ensure high quality of the collected annotations for D2T-Eval.
@@ D.2 Collecting annotations
For the annotation guidelines, we (the authors of this paper) went through an iterative process to establish the annotation guidelines $\mathcal{G}$ and the annotation categories $C$. We started with a preliminary version of the guidelines and annotation categories, drawing inspiration from the guidelines in previous works Kasner & Dušek (2024); Thomson & Reiter (2020). Eventually, we settled on the list of annotation categories described in Appendix E (Table 5).
Our list is an extended and updated version of the error taxonomy of Kasner & Dušek (2024), covering also coherence categories (Incoherent, Repetitive) on top of semantic accuracy.
@@
We provide these additional recommendations in $\mathcal{P}_{\texttt{base}}$, while we omit them in $\mathcal{P}_{\texttt{noguide}}$ (cf. Figures 7 and 8).
@@
Following the established annotation guidelines, we proceeded to collect our own internally-annotated gold data. The purpose of this data is two-fold: (1) to pre-select skilled crowdworkers, and (2) to create a smaller high-quality development set for the model prompting study. Since we also wanted to quantify our own internal IAA, we divided the data into two sets: $\mathcal{D}_{\texttt{dev}}$, containing 84 examples annotated individually by 7 annotators (12 examples per annotator), and $\mathcal{D}_{\texttt{dev}}$, containing 12 examples annotated commonly by each annotator. Our average IAA on $\mathcal{D}_{\texttt{dev}}$ was $F_{1}=0.433$ and $\gamma=0.399$ ([TODO: ] provides the average agreement of crowdworkers and models w.r.t. our annotations on this split).
@@
To gather a pool of skilled annotators, we set up a qualification task on Prolific.com. For the study, we pre-selected workers whose first language is English, with >99% approval rate and more than 100 submissions. The workers were first presented with a detailed tutorial with annotation guidelines and examples of individual errors. After the tutorial, we tested the worker performance on five manually pre-selected examples from $\mathcal{D}_{\texttt{dev}}$. For annotating the data in $\mathcal{D}_{\texttt{test}}$, we selected workers that had $F_{1}$-score higher than 0.5 w.r.t. our ground truth solution to the qualification task. Of the 230 annotators who participated in the qualification task, 50 annotators (21.7%) passed this baseline. We pay all the annotators an average reward of 9.58 GBP / hour, regardless of the qualification outcome.
@@
We invited annotators who passed the qualification task to annotate $\mathcal{D}_{\texttt{test}}$. Out of 50 selected participants, 45 participated in the main task (=90% turnover rate). We presented each annotator with a batch of 32 examples: 25 examples from $\mathcal{D}_{\texttt{test}}$ and 7 remaining examples from $\mathcal{D}_{\texttt{dev}}$ (i.e, the examples from $\mathcal{D}_{\texttt{dev}}$ that we did not use for the qualification task). All the 1200 outputs in $\mathcal{D}_{\texttt{test}}$ were annotated by at least one annotator. Additionally, 475 outputs (39.6%) were annotated by an extra annotator. We payed all the annotators an average reward of 10.70 GBP / hour. We use examples with two annotators to compute the average IAA for D2T-Eval in Section 5.2.151515For other experiments, we use only the outputs from the first annotator as the reference data.
D.3 Annotation Interface
@@
Appendix E Annotation Categories
Tables 5, 6 and 7 show an overview of annotation span categories that we used for our tasks along with their descriptions.
@@ Appendix E Annotation Categories
Appendix F Prompts
Figures 7, 8, 9 and 10 show the prompts for the D2T-Eval that we use for the experiments in Section 5.1.
@@ Appendix G Results
The results for D2T-Eval are included in Sections G.1, G.2, G.3 and G.4:
The results for MT-Eval are included in Sections G.5 and G.6:
The results for Propaganda are included in Section G.7: Tables 20 and 21.
The confusion matrices are included in Section G.8: Figures 14, 15 and 16.
The manual evaluation results are included in Section G.9: Tables 22 and 23.
G.1 D2T-Eval – Main Results
@@
Llama 3.3 0.307 0.132 0.276 0.185 0.388 0.154 0.323 0.169 0.109 0.418
GPT-4o 0.346 0.178 0.300 0.180 0.303 0.179 0.301 0.122 0.130 0.429
Claude 3.7 0.512 0.262 0.395 0.287 0.432 0.274 0.412 0.138 0.203 0.592
DeepS. R1 0.453 0.293 0.493 0.154 0.259 0.202 0.340 0.138 0.185 0.645
o3-mini 0.505 0.351 0.488 0.250 0.347 0.292 0.405 0.113 0.273 0.637
Gem. 2-FT 0.458 0.259 0.434 0.236 0.395 0.247 0.414 0.167 0.209 0.612
@@ G.2 D2T-Evalfootball domain
Llama 3.3 0.512 0.241 0.408 0.295 0.500 0.266 0.449 0.183 0.197 0.455
GPT-4o 0.422 0.271 0.394 0.263 0.382 0.267 0.388 0.121 0.190 0.427
Claude 3.7 0.612 0.330 0.466 0.384 0.542 0.355 0.501 0.146 0.254 0.587
DeepS. R1 0.561 0.414 0.643 0.237 0.368 0.301 0.468 0.167 0.276 0.659
o3-mini 0.610 0.431 0.558 0.344 0.445 0.382 0.495 0.113 0.328 0.645
Gem. 2-FT 0.588 0.387 0.571 0.314 0.464 0.347 0.512 0.165 0.307 0.647

### v2

@@ 6 Discussion
The IAA between LLMs and human annotators is generally moderate. However, we also need to take into account a realistic upper bound: the average IAA between human annotators themselves. In that case, the strongest LLMs have already reached the level of skilled crowdworkers in D2T-Eval. As we also showed, not all human annotations are “gold”, as even qualified crowdworkers make similar amounts of errors as LLMs. If we also factor in financial aspects, LLMs provide better flexibility, shorter response times, and lower costs. In light of this, we argue that the advantages of automating span annotation with LLMs outweigh its issues. LLM-based span annotation will also benefit from future increases in LLM capabilities.
@@
To use an LLM as a span annotator, we strongly recommend providing detailed guidelines describing how to handle all edge cases. In contrast, we note that providing specific examples (i.e., a few-shot setup) does not bring consistent improvements, possibly due to the length and complexity of examples for this task. When using LLMs in a new setup, we recommend validating model’s annotations against examples hand-annotated by experts on a sample of the data. Regarding model choice, we recommend choosing a reasoning model as opposed to an equivalent model without reasoning, since the reasoning models provide more reliable annotations at the cost of slightly higher response times and token count.
@@ 7 Conclusion
We demonstrated that LLMs can serve as span annotators for three span annotation tasks: evaluating data-to-text generation, identifying errors in machine translation, and detecting propaganda in human-written texts. Our experiments show that LLMs achieve moderate agreement with skilled human annotators. The models perform best in D2T-Eval, where they are comparable to qualified crowdworkers. Reasoning models consistently outperform their instruction-tuned counterparts, delivering more accurate annotations and providing more valid explanations for their decisions. Overall, automating span annotation with LLMs offers a promising alternative to human annotation, opening the way for more fine-grained NLG evaluation methods.
@@ Acknowledgments
The human evaluation study was approved by the internal ethics committee of our institution. Our human annotators were hired over Prolific and paid the platform-recommended wage of 9 GBP/hour (adjusted to slightly higher rates to account for real annotation times). Annotators were pre-selected on the basis of their primary language (English). All annotators were shown detailed instructions and explanation of the data types, data sources, and the purpose of the research (see Appendix B.2 for details). The domains were selected so that they do not contain sensitive or potentially offensive content. We do not collect demographic data about participants.
@@ Appendix A Implementation Details
A.1 Open Models
@@
For better reproducibility, we set the seed to 42 and the temperature to 0 for the local models. We do not use these parameters for proprietary models as these parameters are generally not supported.
@@ A.2 Proprietary Models
A.3 Parsing the Output with Reasoning Traces
@@
At the time we performed our experiments, some of the reasoning models either did not support structured output (Gemini 2.0 Flash Thinking) or it would cancel out their ability to perform reasoning (ollama models [^15] https://github.com/ollama/ollama/issues/8529). Therefore, we use the following heuristic: we retrieve the raw answer from the model, strip any parts within the <think></think> tags (if present), and consider the latest valid top-level JSON object as the model’s response.
Appendix B D2T-Eval Details
@@
In this section, we provide more details on the process of collecting data (Section B.1) and annotations (Section B.2) for D2T-Eval.
B.1 Collecting Data
@@
To collect a novel dataset for D2T-Eval that is not subject to data contamination (cf. Section 4.4), we follow the approach of Kasner & Dušek (2024) to collect structured data from publicly available APIs.
@@
The Quintd tool Kasner & Dušek (2024) enables downloading data for five domains: openweather, gsmarena, ice_hockey, our_world_in_data, and wikidata. We selected openweather and gsmarena as the most suitable for the data collection process. Additionally, we collect the new football domain from RapidAPI - API-Football. [^16] Its structured data contain more detailed information about the game than ice_hockey and the sport itself is also more generally well-known.
@@
To obtain output texts, we prompt two state-of-the-art models, Llama 3.3 Grattafiori et al. (2024) and GPT-4o Hurst et al. (2024), and two smaller models, Gemma 2 Team et al. (2024) and Phi-3.5 Abdin et al. (2024) (cf. Section A.1). We prompt the models in a zero-shot setting, asking them to generate a summary of the given data using approximately five sentences. We provide the prompts in Appendix D.
@@
Two of the models (Llama 3.3 and GPT-4o) are also used later to collect annotations, which can reveal self-preference bias Zheng et al. (2023); Wataoka et al. (2024). Since we do not rank the quality of generated outputs, we do not consider it a flaw but rather a ground for future investigation. Moreover, we hypothesize that since span annotations are more grounded than direct assessment, the self-preference bias will not be as prominent.
B.2 Collecting annotations
Here we describe our process of writing annotation guidelines and selecting crowdworkers in order to ensure high quality of the annotations collected for D2T-Eval.
@@ B.2 Collecting annotations
For the annotation guidelines, we (the authors of this paper) went through an iterative process to establish the annotation guidelines $\mathcal{G}$ and the annotation categories $C$. We started with a preliminary version of the guidelines and annotation categories, drawing inspiration from the guidelines in previous works Kasner & Dušek (2024); Thomson & Reiter (2020). Eventually, we settled on the list of annotation categories described in Appendix C (Table 5).
Our list is an extended and updated version of the error taxonomy of Kasner & Dušek (2024), also covering coherence categories (Incoherent, Repetitive) on top of semantic accuracy.
@@
We provide these additional recommendations in $\mathcal{P}_{\texttt{base}}$, while we omit them in $\mathcal{P}_{\texttt{noguide}}$ (cf. Figures 6 and 7).
@@
With the annotation guidelines established, we proceeded to collect our own internally annotated gold data. The purpose of this data is two-fold: (1) to pre-select skilled crowdworkers, and (2) to create a smaller high-quality development set for the model prompting study. Since we also wanted to quantify our own internal IAA, we divided the data into two sets: $\mathcal{D}_{\texttt{dev}}$, containing 84 examples annotated individually by 7 annotators (12 examples per annotator), and $\mathcal{D}_{\texttt{iaa}}$, containing 12 examples annotated commonly by each annotator. Our average IAA on $\mathcal{D}_{\texttt{iaa}}$ was $F_{1}=0.444$ and $\gamma=0.399$.
@@
To gather a pool of skilled annotators, we set up a qualification task on Prolific.com. For the study, we pre-selected workers whose first language is English, with >99% approval rate and more than 100 submissions. The workers were first presented with a detailed tutorial with annotation guidelines and examples of individual errors. After the tutorial, we tested the worker performance on five manually pre-selected examples from $\mathcal{D}_{\texttt{dev}}$. For annotating the data in $\mathcal{D}_{\texttt{test}}$, we selected workers that had $F_{1}$ score higher than 0.5 w.r.t. our ground-truth solution to the qualification task. Of the 230 annotators who participated in the qualification task, 50 annotators (21.7%) passed this baseline. We pay all the annotators an average reward of 9.58 GBP / hour, regardless of the qualification outcome.
@@
We invited annotators who passed the qualification task to annotate $\mathcal{D}_{\texttt{test}}$. Out of 50 selected participants, 45 participated in the main task (=90% turnover rate). We presented each annotator with a batch of 32 examples: 25 examples from $\mathcal{D}_{\texttt{test}}$ and 7 remaining examples from $\mathcal{D}_{\texttt{dev}}$ (that is, the examples from $\mathcal{D}_{\texttt{dev}}$ that we did not use for the qualification task). All 1200 outputs in $\mathcal{D}_{\texttt{test}}$ were annotated by at least one annotator. Furthermore, 475 outputs (39.6%) were annotated by an additional annotator. We payed all the annotators an average reward of 10.70 GBP / hour. We use examples with two annotators to compute the average IAA for D2T-Eval in Section 5.2.171717For other experiments, we use only the outputs from the first annotator as the reference data.
B.3 Annotation Interface
@@
Appendix C Annotation Categories
Tables 5, 6 and 7 show an overview of the annotation span categories that we used for our tasks along with their descriptions.
@@ Appendix C Annotation Categories
Appendix D Prompts
Figures 6, 7, 8 and 9 show the prompts for the D2T-Eval that we use for the experiments in Section 5.1.
@@ Appendix E Results
The results for D2T-Eval are included in Sections E.1, E.2, E.3 and E.4:
The results for MT-Eval are included in Sections E.5 and E.6:
The results for Propaganda are included in Section E.7: Tables 20 and 21.
The confusion matrices are included in Section E.8: Figures 13, 14 and 15.
The manual evaluation results are included in Section E.9: Tables 22 and 23.
E.1 D2T-Eval – Main Results
@@
Llama 3.3 0.307 0.176 0.365 0.187 0.388 0.181 0.377 0.196 0.109 0.418
GPT-4o 0.346 0.233 0.391 0.184 0.308 0.206 0.345 0.139 0.130 0.429
Claude 3.7 0.512 0.294 0.442 0.304 0.457 0.299 0.449 0.150 0.203 0.592
DeepS. R1 0.453 0.317 0.532 0.185 0.310 0.233 0.392 0.159 0.185 0.645
o3-mini 0.505 0.392 0.542 0.285 0.395 0.330 0.457 0.127 0.273 0.637
Gem. 2-FT 0.458 0.293 0.488 0.263 0.438 0.277 0.462 0.185 0.209 0.612
@@ E.2 D2T-Evalfootball domain
Llama 3.3 0.512 0.288 0.484 0.296 0.498 0.292 0.491 0.199 0.197 0.455
GPT-4o 0.422 0.347 0.502 0.268 0.388 0.302 0.438 0.136 0.190 0.427
Claude 3.7 0.612 0.360 0.511 0.391 0.555 0.375 0.532 0.157 0.254 0.587
DeepS. R1 0.561 0.431 0.670 0.275 0.428 0.336 0.522 0.186 0.276 0.659
o3-mini 0.610 0.470 0.608 0.384 0.496 0.422 0.547 0.125 0.328 0.645
Gem. 2-FT 0.588 0.419 0.617 0.347 0.512 0.379 0.560 0.181 0.307 0.647